OpenStudy (anonymous):

The area of a particular rectangle is 12 times the area of a certain square, and the width of the rectangle is three times the length of a side of the square. The perimeter of the rectangle is 30 units greater than the perimeter of the square. Find the dimensions of both the rectangle and the square.

OpenStudy (nikato):

Area of rectangle=wL. Square=s^s

OpenStudy (nikato):

So wL=12s^2 w=3s 2w+2L=30+4s

OpenStudy (anonymous):

Okay I'm getting it.

OpenStudy (nikato):

Ok, substitute the second equation into the first. And u should het something for L in terms of s

OpenStudy (anonymous):

gotcha

OpenStudy (nikato):

So what did u get?

OpenStudy (anonymous):

12s^2/3s

OpenStudy (nikato):

And u can simplify that to 4s

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

wait....It simplifies to 4s^3.....

OpenStudy (nikato):

So u have w=3s and L=4s Substitute those values for w and L into the last equation and u should be able to find s, which is one side of the square, which is also all 4 sides

OpenStudy (nikato):

No, it just 4s

OpenStudy (nikato):

Were u able to find ur answers?

OpenStudy (anonymous):

okay. Yeah thanks

OpenStudy (nikato):

Ok. No problem. But just checking. Square=3x3 Rectangle =9x12

OpenStudy (anonymous):

YUP! Thanks